# Boosting Graph Foundation Model from Structural Perspective

### v1

@@ 4.2 Unifying Structure from Diverse Domains
where $W_{3}\in\mathbb{R}^{d\times d}$ is the learnable parameters and $\beta$ is a hyper-parameter. Instead of relying on the original structure, we introduce super nodes and connect them to all nodes within their neighborhood by virtual edges. This new structure allows for effective information aggregation while unifying structure space across domains, thus boosting the model generalizability.
4.3 Pre-training
To optimize the pre-trained model, we leverage a contrastive loss function $\mathcal{L}_{pre}$. The super node $p$ aggregates information from its neighbor node $u\in\mathcal{N}(u)$ based on learned weights $a_{pu}$. Intuitively, neighbors that are more similar to the super node are more important. Therefore, for each encoded super node $\hat{h}(p_{ij})$, the original super node $h(p_{ij})$ before encoding is treated as a positive sample. Meanwhile, encoded super nodes from different classes are selected as negatives. We formulate the loss as the normalized temperature-scaled cross-entropy loss [1] as follows:
$\mathcal{L}_{pre}=-\frac{1}{C}\sum_{j}^{C}\frac{1}{n}\sum_{i}^{n}log\frac{exp(sim(\hat{h}(p_{ij}),h(p_{ij})/\tau)}{\sum_{q=1,q\neq j}^{C}(exp(sim(\hat{h}(p_{ij}),\hat{h}(p_{iq}))/\tau))}$ (6)
where $C$ is the number of classes, $n$ is the number of instances, $\tau$ is a temperature hyper-parameter and sim(·) is implemented with cosine similarity.
Similarity Matching & Zero-shot learning. We employ similarity matching to select the final representation for each input graph/node. Let $z_{i}$ denote the final representation of $i$-th input graph/node, it can be computed as:
$z_{i}=\hat{h}(p_{iw}),$ (7)
where $w=\arg\max_{j}(sim(\hat{h}(p_{ij}),c(l_{j})))$, $j\in\left[1,C\right]$. Intuitively, super nodes that effectively aggregate information within their neighborhoods are most similar to their corresponding class labels. Therefore, the index $w$ based on the maximum similarity calculation is also the predicted class for node or graph classification tasks. Then, we define the prediction for the $i$-th graph/node as:
$\hat{y}_{i}=\arg\max_{j}(sim(\hat{h}(p_{ij}),c(l_{j}))),j\in\left[1,C\right].$ (8)
Since predictions based on similarity matching do not rely on labels, and the pre-trained model can generalize to unseen data. We treat the similarity matching process as zero-shot learning. It is worth noting that when evaluating the model performance, we keep the parameters of the pre-trained model frozen.
For link prediction tasks in the zero-shot setting, we compare the similarity of input node pairs. Pairs with similarity greater than a threshold $T$ are classified as having an edge; otherwise, they are considered not to have an edge. Let $\hat{e}_{ij}$ denote the link prediction result between the pair of nodes $(v_{i},v_{j})$. It can be computed as follows:
$\hat{e}_{ij}=\begin{cases}1,\text{if}~{}sim(z_{i},z_{j})>T\\ 0,\text{otherwise}\end{cases}$ (9)
@@ 4.4 Tuning for Downstream Tasks
Next, we show how we freeze the pre-trained model and adapt it to different downstream tasks on graphs. BooG uses MLP as the classifier to predict the class of instances. Formally, the final prediction for $i$-th input graph/node is given by:
$\hat{y_{i}}=\arg\max(MLP(z_{i})).$ (10)
In link prediction tasks, we standardize it as a binary classification task and employ the concatenated representation of node pairs as the input to the MLP for prediction. In both supervised and few-shot learning scenarios, we employ cross-entropy loss to train the MLP for all tasks.
@@ 4.5 Time Complexity
We analyze the time complexity of BooG in this section. For BooG, we denote the maximum sequence length of node textual feature as $L$ and the number of instances as $n$. For an input graph/node, the complexity is $O(L^{2}d+Ld^{2})$, where $d$ is the dimensionality of the embeddings. The time complexity is $O(n*(L^{2}d+Ld^{2}))$ for processing $n$ instances. The encoder’s time complexity primarily depends on the feature dimensionality and the structure of the graph. The time complexity of the attention mechanism can be expressed as $O(2d^{2}n+2dn)$. Next, the complexity of message aggregation is $O(\bar{k}dn+d^{2}n)$, where $\bar{k}$ is the average number of neighbors per node. The overall time complexity of encoder is $O(k_{1}d^{2}n+k_{2}dn)$, where $k_{1},k_{2}$ are the coefficients. For similarity matching, the time complexity is $O(dnC)$, where $C$ is the number of classes. Overall, our method BooG has a linear time complexity w.r.t. $n$.
5 Experiments
In this section, we evaluate the generalizability of BooG on both node-level and graph-level tasks.
5.1 Experimental Settings
Datasets. We conduct experiments on six benchmark datasets, which include 4 node-level graph datasets (Cora, Pubmed, ogbn-arxiv, Wiki-CS) and 2 graph-level datasets (PCBA, HIV). Due to the space limitation, we move details on datasets (Sec. A), experimental setup (Sec. B) and hyper-parameter sensitivity analysis (Sec. C) to Appendix.
Baselines. We compare BooG with six baselines, which can be categorized into four types. (1) MLP utilizes a multi-layer perceptron to extract deep features individually for each node. (2) Graph Neural Networks: GAT [25] computes the hidden representations of each node in the graph by first learning the importance of its neighbors and then aggregating information from them. GIN [29] enhances the representation power of GNNs by employing a distinct graph encoding method that emphasizes the discrimination of non-isomorphic structures. (3) Graph Pre-training Models: GraphCL [30] utilizes pre-training of graph models through the application of a self-discriminative contrastive learning task on learned node embeddings. DGI [26] introduces a self-supervised pre-training task that aims to maximize the mutual information between the local node view and the global graph view. (4) Graph foundation models: OFA [14] utilizes a pre-trained LM to align feature spaces of graphs from different domains, enabling supervised training across all graphs.
5.2 Performance Results
We next evaluate the model generalizability across various learning scenarios and downstream tasks. We employ TAGs for all methods to unify all input data. Specifically, we employ the pre-trained language models (LMs) to encode instance features as input for models. BooG and all other self-supervised learning baselines are first pre-trained on the citation network Cora. In the supervised learning setting, the parameters of pre-trained models are frozen and the MLP is trained in a supervised manner. Other baselines are trained directly on the target datasets. In the few-shot learning setting, we freeze the parameters of the pre-trained methods and train their MLPs using a $N$-way $K$-shot task. For other methods, we train them using the $N$-way $K$-shot task, i.e., they are directly trained on target datasets with labels. To accomplish this, we construct a support set by randomly selecting $K$ examples per class from the training split. For zero-shot learning, different from the above two scenarios, OFA is first pre-trained on the citation network Cora. For GraphCL and DGI, predictions are generated by comparing the similarity between the instance representations derived from their models and the class label representations encoded by the LM. Finally, we evaluate all methods on the test sets of the target datasets.
Performance analysis. Table 1, Table 2 and Table 3 report the results in supervised learning, few-shot learning, and zero-shot learning, respectively. From these tables, we observe that: (1) BooG demonstrates the best results in most cases. (2) The GNN models GAT and GIN perform better in specific domains in supervised learning and few-shot learning. For example, GAT excels on citation network datasets, while GIN is more suited to molecular graphs. However, these methods perform poorly in other domains because they are typically designed for a single graph domain, especially in few-shot learning. (3) The two self-supervised learning methods, GraphCL and DGI, exhibit poor domain generalization capabilities in few-shot learning and zero-shot learning, since they only use the raw graph structure. (4) OFA demonstrates a certain degree of domain generalization capabilities in few-shot learning, highlighting the effectiveness of in-context learning. However, BooG outperforms OFA in most cases because it constructs unified structural characteristics for graphs from different domains, thereby enhancing the model generalizability. Overall, BooG shows strong generalization ability in all scenarios, achieving results comparable to or even better than baseline methods, which proves its effectiveness.
Datasets and tasks MLP GAT GIN OFA GraphCL DGI BooG
[…]
Datasets and tasks MLP GAT GIN OFA GraphCL DGI BooG
[…]
Datasets and tasks Cora Pubmed ogbn-arxiv Wiki-CS Cora-link PCBA HIV
[…]
DGI 38.36 28.46 24.11 23.76 35.85 21.02 25.11
[…]
Figure 4: t-SNE visualization on Cora.
Figure 5: Ablation study.
5.3 Visualization
We further evaluate the expressive power of BooG. We visualize node representations before and after pre-training on Cora with the t-SNE technique [24]. Colors are used to distinguish different node classes. Before pre-training, the node representations are encoded by the LM. As shown in Figure 5, the node representations of different classes are more clearly separated after pre-training. This indicates that BooG can learn more expressive representations.
5.4 Ablation Study
The ablation study is done to understand the importance of main components of BooG in supervised learning and 5-shot learning scenarios on Cora and HIV. The model is pre-trained on Cora. We first remove the class labels from the super nodes and call this variant BooG-nc (no class label). Accordingly, in the pre-training task, we use the anchor node before aggregation as the positive sample and treat other instances of dataset as negative samples. Secondly, we remove the anchor nodes from the super nodes and treat the anchor node as a member of the contextual neighborhood. We call this variant BooG-na (no anchor node). Further, we remove the similarity matching process and concatenate the representations based on different class labels as the final representation of the input to the MLP. We call this variant BooG-ns (no similarity matching). We compare BooG with these three variants, and the results are presented in Figure 5. Our findings show that BooG outperforms all the variants on the two datasets. Further, the decline in performance exhibited by both BooG-nc and BooG-na highlights the importance of super nodes. In other words, the super nodes significantly enhance model generalizability, especially in graph domain transfer.
6 Conclusion
In this paper, we boost graph foundation model from structural perspective and propose BooG. The model constructs virtual super nodes to unify structural characteristics of graph data from different domains. Specifically, the virtual super nodes fuse the information of both anchor nodes and class labels, where each anchor node captures the information of a node instance or a graph instance to be classified. Instead of utilizing the raw graph structure, we introduce super nodes and connect them to all the nodes within the neighborhood of an anchor by virtual edges. This new linkage allows for effective information aggregation while disregarding the difference in structural connectivity across domains. In this way, cross-domain model generalization can be boosted from structural perspective. Additionally, we propose a novel pre-training objective based on contrastive learning, which learns more expressive representations for graph data and generalizes effectively to different domains and downstream tasks. Our extensive experiments and ablation studies validate the effectiveness of BooG.
Appendix A Datasets
In this section, we introduce the datasets used in the experiments. The statistics of datasets are listed in Table 4, and we describe each dataset in detail below:
Cora is a homophilous graphs commonly used as benchmarks. It consists of 2708 scientific publications classified into one of 7 classes. In this dataset, each node represents a scientific paper and each edge represents a citation. We collect raw text from [5].
Pubmed is a homophilous graphs commonly used as benchmarks. It consists of 19,717 scientific publications from PubMed database pertaining to diabetes classified into one of 3 classes. We collect raw text from [5].
ogbn-arXiv is an academic citation network graph dataset, consisting of papers from arXiv and their citation relationships. Each node represents a paper, and edges represent the citation relationships between papers. The task is to predict the 40 subject areas of arXiv CS papers. We collect raw text from [8].
Wiki-CS is an Internet link network where each node represents a Wikipedia page and each edge represents a reference link. Each node’s label corresponds to the category of the entry. We collect raw text from [14].
HIV is a subset of the PubChem BioAssay dataset, focusing on predicting the activity of chemical molecules against the HIV virus. The task in the HIV dataset involves predicting the activity of chemical molecules against the HIV virus. We collect raw text from [14].
For Cora and PubMed, we measure the performance of all models on the test sets over 10 random splits as suggested in [20] and report the average accuracy. For ogbn-arxiv, we follow the official splits [8]. Following the experimental procedure suggested by OGB, we repeat each experiment for 10 times with random seeds and report the average accuracy. For Wiki-CS, we follow the official splits [19] with 20 different training splits and report the average accuracy. For PCBA and HIV, we follow the official splits [32], repeat each experiment for 10 times with random seeds and report the average accuracy. The statistics and details of these datasets can be found in Table 4.
@@ Appendix A Datasets
Appendix B Experimental Setup
We implement BooG by PyTorch and conducted the experiments with one A100 GPU. The model is optimized by Adam [10]. We employ TAGs for all methods to unify all input data. Specifically, we employ the pre-trained language models (LMs) to encode instance features as input for models. BooG and all other self-supervised learning baselines are first pre-trained on the citation network Cora. For other self-supervised learning baseline methods, we train a linear classifier on the top of the embeddings derived from the frozen model. Then, we evaluate the model performance on the test sets of all six datasets with different tasks. We perform a grid search to tune hyper-parameters based on the validation set. Details on the search space of these hyper-parameters can be found in Table 5.

### v2

@@ 4.2 Unifying Graph Structures
where $W_{3}\in\mathbb{R}^{d\times d}$ is a learnable parameters and $\beta$ is a hyper-parameter.
4.3 Self-Supervised Learning Paradigm
Pre-training. We pre-train BooGvia a contrastive learning objective to enhance its ability to distinguish between different class hypotheses and improve representation generalizability.
For each anchor node $s_{i}$, we generate $C$ super nodes $\{p_{ij}\}_{j=1}^{C}$ corresponding to all possible class labels. These super nodes incorporate both the anchor node’s information and individual class semantics through Equation (5). After neighborhood aggregation, we obtain the updated embedding $\hat{h}(p_{ij})$ for each super node $p_{ij}$.
To guide the model toward learning class-discriminative representations, we adopt a normalized temperature-scaled cross-entropy loss [63], formulated as:
$\begin{split}&\mathcal{L}_{pre}=\\ &-\frac{1}{C}\sum_{j}^{C}\frac{1}{n}\sum_{i}^{n}log\frac{\exp(\text{sim}(\hat{h}(p_{ij}),h(p_{ij}))/\tau)}{\sum_{q=1,q\neq j}^{C}(\exp(\text{sim}(\hat{h}(p_{ij}),\hat{h}(p_{iq}))/\tau))}.\end{split}$ (8)
Here, $n$ denotes the number of training instances, $C$ is the number of classes, $\tau$ is a temperature hyper-parameter, and $\text{sim}(\cdot)$ computes cosine similarity. Notably, while the loss function considers the number of classes $C$ in the dataset, it does not necessitate concrete sample-label pairs. Our pre-training strategy leverages self-supervised learning via contrastive learning.
In this framework, the original embedding $h(p_{ij})$ of super node $p_{ij}$—formed by the combination of the anchor node $s_{i}$ and class label $l_{j}$—is treated as the positive sample for its updated embedding $\hat{h}(p_{ij})$. Meanwhile, embeddings $\hat{h}(p_{iq})$ with $q\neq j$ serve as negative samples for the same anchor node $s_{i}$, as they are associated with different class labels. This contrastive objective explicitly drives the model to align the aggregated representation $\hat{h}(p_{ij})$ closer to its corresponding class label (positive pair) while pushing it away from other class hypotheses (negative pairs). This alignment forces the model to learn more discriminative features that capture class semantics, thereby enhancing its expressive power and its ability to generalize across tasks and domains.
Similarity Matching. After pre-training, each anchor node $s_{i}$ produces $C$ super node embeddings, one for each class label. To select the most appropriate representation for downstream tasks, we apply similarity matching:
$z_{i}=\hat{h}(p_{iw}),\quad\text{where}\quad w=\arg\max_{j\in[1,C]}\text{sim}(\hat{h}(p_{ij}),c(l_{j})),$ (9)
where $c(l_{j})$ is the embedding of the $j$-th class label. This step ensures that the final embedding $z_{i}$ for $s_{i}$ is selected based on the class label most semantically aligned with the aggregated neighborhood information.
@@ 4.4 Tuning for Downstream Tasks
Zero-shot Scenarios. Since the predictions provided by similarity matching do not depend on labels, the pre-trained model can generalize to unseen data, allowing us to treat this process as zero-shot learning. It is worth noting that the parameters of the pre-trained model are kept frozen during evaluation under all learning scenarios. For node/graph classification tasks, the prediction for the $i$-th node/graph is defined as:
$\hat{y}_{i}=\arg\max_{j}(\text{sim}(\hat{h}(p_{ij}),c(l_{j}))),j\in\left[1,C\right].$ (10)
For link prediction tasks in the zero-shot setting, we compare the similarity of input node pairs. Pairs with a similarity exceeding a threshold $T$ are classified as having an edge; otherwise, they are considered not to have an edge. The link prediction result $\hat{e}_{ij}$ between the pair of nodes $(v_{i},v_{j})$ is computed as follows:
$\hat{e}_{ij}=\begin{cases}1,\text{if}~\text{sim}(z_{i},z_{j})>T\\ 0,\text{otherwise}\end{cases}.$ (11)
Supervised & Few-shot Scenarios. BooG uses MLP as the classifier to predict the class of instances. Formally, the final prediction for $i$-th input node/graph is given by:
$\hat{y_{i}}=\arg\max(\text{MLP}(z_{i})).$ (12)
For link prediction tasks, the problem is framed as a binary classification task, where the concatenated representation of node pairs is used as input to the MLP for prediction. In both supervised and few-shot learning scenarios, we employ cross-entropy loss to train the MLP across all tasks.
@@ 4.5 Time Complexity
We analyze the time complexity of BooG in this section. For BooG, we denote the maximum sequence length of node textual feature as $L$ and the number of instances as $n$. For an input graph/node, the complexity is $O(L^{2}d+Ld^{2})$, where $d$ is the dimensionality of the embeddings. The time complexity is $O(n*(L^{2}d+Ld^{2}))$ for processing $n$ instances. The encoder’s time complexity primarily depends on the feature dimensionality and the structure of the graph. The time complexity of the attention mechanism can be expressed as $O(2d^{2}n+2dn)$. Next, the complexity of message aggregation is $O(\bar{k}dn+d^{2}n)$, where $\bar{k}$ is the average number of neighbors per node. The overall time complexity of encoder is $O(k_{1}d^{2}n+k_{2}dn)$, where $k_{1},k_{2}$ are the coefficients. For similarity matching, the time complexity is $O(dnC)$, where $C$ is the number of classes. Overall, BooG has a linear time complexity w.r.t. $n$.
@@
4.6 Why BooG Works?
BooG boosts model’s cross-domain and cross-task generalization ability from two key aspects. First, BooG unifies graph attributes by encoding node texts of TAGs and class label descriptions into a shared space using a pre-trained LM, thereby equipping the model to handle graph data from diverse domains. More critically, BooG unifies graph structures by introducing the concept of sub-graphs and super nodes. Sub-graphs offer a consistent task template across all levels of tasks, consisting of anchor nodes and their neighbor nodes. The super nodes, along with virtual edges, establish a standardized aggregation mechanism that fuses rich information from neighborhoods and associated class labels, accommodating graph structural characteristics inherent to different domains. Consequently, these designs integrate a supernode-centered graph structure that encapsulates label preferences, making BooG compatible with different graph domains and task levels.
Recent studies on graph domain generalization include UniGraph [48] and OFA [46]. UniGraph proposes a cascade architecture integrating LM and GNN, utilizing a self-supervised objective based on Masked Graph Modeling. Additionally, it incorporates graph instruction fine-tuning through LLMs to enable zero-shot prediction. BooG freezes LM throughout the process and fundamentally resorts to the unification of graph attributes and structures as the primary vehicle for generalization. In addition, while OFA leverages an LM to encode node features across domains and introduces the nodes-of-interest (NOI) subgraph and NOI prompt node to unify different graph tasks, BooG’s super nodes are designed to bind anchor nodes directly with class labels, as expressed in Eq.(5), which aligns them explicitly within the same latent space, ensuring a more cohesive integration between node features and task-specific labels.
5 Experiments
5.1 Experimental Settings
Datasets. We first introduce the datasets used in the experiments. The statistics of these datasets are summarized in Table 1, and details of each dataset are given below:
Cora is a classic graph dataset widely used in research on GNNs and node classification tasks. It consists of 2708 scientific publications classified into one of 7 classes. In this dataset, each node represents a scientific paper and each edge represents a citation. The label of nodes indicates the research field of the paper. We collect raw text from [64].
Pubmed is another widely used dataset for graph-based machine learning, particularly in the context of node classification tasks with GNNs. It consists of 19,717 scientific publications classified into one of 3 classes. Similar to the Cora dataset, it is a citation network of scientific papers, but it consists of biomedical papers indexed in PubMed, which is a free search engine accessing primarily life sciences and biomedical literature. We collect raw text from [64].
ogbn-arXiv is part of the Open Graph Benchmark (OGB), a collection of large-scale graph datasets designed to evaluate machine learning algorithms, particularly for graph-based learning tasks. It is focused on node classification in a citation network of academic papers, specifically from the arXiv repository, which is a preprint server for scientific papers in fields such as physics, mathematics, computer science, and more. Each node represents a paper, and edges represent the citation relationships between papers. The task is to predict the 40 subject areas of arXiv CS papers. We collect raw text from [65].
Wiki-CS is another graph dataset in the OGB collection, specifically designed for node classification tasks within the context of social networks. It is based on Wikipedia articles and represents a graph of scientific papers. Each node represents a Wikipedia page and each edge represents a reference link. Each node’s label corresponds to the category of the entry. We collect raw text from [46].
Ele-Fashion[66] is a node classification dataset derived from the Amazon-Fashion dataset. Each node represents a fashion product, and links indicate user co-purchasing behavior. Textual features are product titles. We collect raw text from[66]. The task is to predict product categories, simulating recommendation scenarios.
HIV is a collection of data used to study issues related to the treatment of HIV (Human Immunodeficiency Virus). It is commonly used in machine learning and data analysis tasks such as drug discovery, drug-virus interaction prediction, and drug resistance prediction. It is a subset of the PubChem BioAssay dataset. The task in the HIV dataset involves predicting the activity of chemical molecules against the HIV virus. We collect raw text from [46].
For Cora and PubMed, we measure the performance of all models on the test sets over 10 random splits as suggested in [67] and report the average accuracy. For ogbn-arxiv, we follow the official splits [65]. Following the experimental procedure suggested by OGB, we repeat each experiment for 10 times with random seeds and report the average accuracy. For Wiki-CS, we follow the official splits [68] with 20 different training splits and report the average accuracy. For PCBA and HIV, we follow the official splits [69], repeat each experiment for 10 times with random seeds and report the average accuracy.